A sportswriter makes his pre-season picks for the top ten teams finish. If there are forty teams, how many different lists could there be?
Answer by greenestamps(13198)



It's not clear from the statement of the problem whether the list is ordered 1st through 10th, or if it is an unordered list of the top 10 teams.

It is probably an ordered list, which means the problem is about permutations. In that case the number of possible different lists is

P(40,10) = 40*39*38*...*32*31 = (large number)

If the list is not ordered, then the problem is about combinations; in that case the number of possible different lists is

C(40,10) = (40*39*38*...*32*31)/(10*9*8*...*2*1) = 847660528